OpenStudy (anonymous):

Let f(x, y) = x2 - xy. Determine a simplified expression for f(x, x-y).

OpenStudy (anonymous):

And that's x^2

OpenStudy (bahrom7893):

f(x,y) = x^2 - xy f(x, x-y) = x^2 - x(x-y) = x^2 - x^2 +xy = xy

OpenStudy (rulnick):

f(x, x-y) = x^2 - x(x-y) = x^2 - x^2 + xy = xy. Yup, bahrom has it!

OpenStudy (anonymous):

Okay so just replace x with x-y? Awesome! Thanks guys.

OpenStudy (rulnick):

Replace y with x-y.

OpenStudy (anonymous):

Oh, yeah. Gotcha. Thanks so much!